# T HREAD : Thinking Deeper with Recursive Spawning

### v1

@@ Results.
Thread significantly outperforms all prior methods (Tables 1 and 2), including those that require the agent to have access to external memory consisting of past experiences with the task. With GPT-4, Thread achieves a combined success rate of 98.5% with 100% success in 4 of the 6 individual tasks. With GPT-3.5, Thread shows a combined success rate of 95.5%, outperforming all methods by over 9% absolute points, and achieves over 90% accuracy across all individual tasks. With smaller models, including Llama-3-8b and CodeLlama-7b, Thread improves upon prior methods by 40% to 55% absolute points.
Finally, Table 3 shows the results when testing Thread with task-general prompting. To implement the task-general prompt, we split the prompt into one set of examples for the main thread and a second set of examples for all other threads (with the same sets of examples used for all tasks). We see that GPT-3.5 achieves the same combined success rate with the task-general prompting as it does with the task-specific prompting. Further, while the performance of Llama-3-8b and CodeLlama-7b degrades with task-general prompting, the performance is still a significant improvement over task-specific prompting with prior methods (Table 2) for these models.
4.2 TextCraft
TextCraft is a text-based environment inspired by the the crafting component of Minecraft [25]. The tasks involve building Minecraft items with crafting commands using available resources from the environment. Following the work of Prasad et al., [25], we evaluate Thread on the test set containing 200 samples. We implement Thread using a single few-shot prompt and compare its performance to previous methods used for TextCraft [25, 33, 27, 37].
Results.
As shown in Table 4, Thread outperforms prior methods by at least 20% absolute points with GPT-3.5 and at least 40% absolute points with Llama-3-8b and CodeLlama-7b. Further, Llama-3-8b with Thread outperforms GPT-3.5 with all prior methods and CodeLlama-7b with Thread outperforms GPT-3.5 with all prior methods except TDAG [33].
Table 4: TextCraft success rate (%). *As reported in Prasad et al., [25].
@@
Reflexion* [27] 32.0 - - -
ReAct [37] 20.5 12.5 8.0 10.5
ADaPT [25] 52.5 23.5 12.0 18.0
TDAG [33] 73.5 48.5 14.0 31.0
@@
4.3 WebShop
WebShop is an online shopping environment with over a million real-world products [36]. The benchmark requires the model to interact with the website (e.g., search, click buttons) to purchase a product based on specifications provided by a user. The search results and product pages contain noisy data with the product title, descriptions, and options. Successful purchase of a product involves selecting the correct item and, if necessary, selecting the correct buttons representing the different item options. The evaluation metrics include the success rate, which is the percentage of products that were purchased with full success (i.e., the purchased item contains all desired attributes), and the score, which is the average percentage of desired attributes covered by the purchased items. Following prior work [27, 25, 33, 40], we evaluate Thread on a test set of 100 instructions. We implement Thread with a few-shot prompt.
@@ 4.3 WebShop
Requires ext. mem. Method GPT-3.5 Llama-3-8b Llama-2-7b CodeLlama-7b
@@
Yes Reflexion [27] 38 64.4 32 59.8 8 19.7 17 57.3
LATS [40] 40 76.0 34 61.5 12 30.1 21 60.7
RAP* [15] 48 76.1 - - - - - -
AutoGuide* [5] 46 73.4 - - - - - -
No ReAct [37] 37 59.5 31 54.1 10 18.5 17 49.2
ADaPT [25] 44 60.0 35 58.2 13 28.7 21 56.4
TDAG [33] 45 64.5 37 63.5 14 29.8 23 58.5
@@ Results.
Table 5 shows the success rate and score for each method. Similar to above, we separate the results based on whether the method requires the model to have access to external memory. With GPT-3.5, Thread outperforms other prompt-only methods by an absolute 4% in success rate and over 10% in score and outperforms RAP [15] by 1% in success rate (with a similar score). Thread achieves an absolute 10%, or greater, improvement in performance with Llama-3-8b and CodeLlama-7b. In addition, Llama-3-8b with Thread achieves a higher success rate than GPT-3.5 with all prior methods, with the exception of RAP.
@@ 4.4 DataCommons QA
DataCommons QA is a benchmark consisting of questions that can be answered using data provided by Google DataCommons [9] https://datacommons.org. The questions range from comparing statistics in different locations to making predictions regarding future trends. Figure 5 of the appendix shows an example with the question “From 2015 to 2021, was the rate of asthma increasing faster in Boston or LA?”. The benchmark includes ground-truth text-based answers (e.g., “Boston”), along with figures that support each answer, according to the data that is available in DataCommons. In this work, we evaluate the models based on the text-based answers, leaving the evaluation of figure generation for future work. The test set includes a total of 140 questions. We implement Thread with a few-shot prompt and compare its performance to three baselines: Reflexion [27], Natural Language Embedded Programs (NLEP) [38], and NLEP+ReAct, an extension of NLEP that allows the model to evaluate intermediate outputs of its analysis as it answers the question.
@@ Results.
Table 6 shows the accuracy of each method on DataCommons QA. Thread outperforms prior methods by over 35% absolute points with GPT-3.5, Llama-3-8b, and CodeLlama-7b. Both Llama-3-8b and CodeLlama-7b with Thread outperform GPT-3.5 with prior methods.
4.5 MIMIC-III ICU QA
The MIMIC-III ICU QA benchmark consists of patient-focused questions based on clinical time-series data made available by MIMIC-III [14]. The benchmark reflects a setting in which a healthcare provider can ask natural language questions about patients in the intensive care unit (ICU) (e.g., evaluating correlations between clinical variables or comparing a patient’s values to other patients matching a specific clinical profile) and receive an answer from the language model based on the relevant patient data. Further details regarding the benchmark and its creation are included in the appendix. The test set includes 160 questions. Similar to above, we implement Thread with a few-shot prompt and compare its performance relative to Reflexion [27], NLEP [38], and NLEP+ReAct.
Results.
As shown in Table 7, Thread outperforms prior methods by over 25% absolute points across all models except Llama-2-7b. All methods show the lowest performance with Llama-2-7b, which is consistent across all benchmarks.
@@
Reflexion [27] 37.9 24.3 10.7 20.7
NLEP [38] 28.6 21.4 11.4 19.3
@@
Reflexion [27] 38.1 19.4 8.8 16.9
NLEP [38] 35.6 17.5 8.1 15.6
@@ 5 Related work
The Thread framework relates to prior approaches for dynamically adapting work, or computational steps, used during model generation. Previous approaches for dynamically adapting computation based on the given problem require special training [7, 22] or novel model architectures [8, 2, 3]. Thread provides a more general and flexible framework that, as we show, can be applied without modifications to the underlying model architecture and without further training. In addition, unlike methods that allow the model to adapt the amount of internal computations [7, 8, 2, 3], the intermediate work used to supplement model generation with Thread can involve work that extends beyond internal thinking, such as retrieving information or interacting with an external environment. In addition, Thread allows for greater interpretability, since all of the intermediate work is performed by generating meaningful tokens.
@@
In this paper, we apply Thread in a setting where it improves the ability of LLMs to decompose a given problem into simpler sub-problems. There has been significant prior work involving the decomposition of problems with neural modular modeling architectures [1, 31, 20, 13, 11, 24, 16]. Later work has used fine-tuning or in-context learning with modern LLMs for decomposition with multi-step task completion [17, 32], arithmetic and mathematical reasoning [6, 18], and program synthesis [21, 23, 39]. To handle complex tasks, a more recent approach has improved upon these methods by using a planner to further decompose a task when a failure is encountered [25], which was further extended to a multi-agent setting [33]. Compared to these approaches, Thread has the advantage of allowing the model to adapt its decision-making in real time as it receives feedback from each step. In addition, Thread provides a more unified framework compared to these methods, which require separate prompting for separate planner and executor modules. Thread instead can be implemented with the same few-shot prompt used for every thread at every step of model generation.
6 Limitations
The implementation of Thread in this paper does not involve any explicit mechanisms for error handling and instead relies on the model’s inherent ability to troubleshoot when issues arise. An advantage of Thread is that individual threads can fail without directly affecting other threads. However, the context necessary to self-correct may be lost if error reporting is not handled appropriately. More work is needed to develop robust error detection and recovery mechanisms to ensure that the information necessary for self-correction is preserved and utilized effectively. In addition, our implementation of Thread involves limited communication between the parent and child threads, where the context for the child thread is based on the last line of the parent’s token sequence. This can result in the child missing information that could help with its work, leading to less efficient or effective generation. Overall, more work is need to improve the propagation of information, as defined by the functions $\phi$ and $\psi$, between parent and child threads.
7 Broader impact
In this work, we show how Thread can be applied to improve LLM task completion and question answering. LLMs are being increasingly deployed to autonomously interact with external environments and humans. By improving this capacity, our work has the potential for amplifying risk associated with automated decision-making or facilitate harmful use of LLMs. Addressing these risks requires careful consideration of ethical guidelines, robustness checks, and transparency in algorithm deployment to ensure that advancements in LLMs contribute positively to societal welfare.
@@ 8 Conclusion
We introduce Thread, a general framework in which model generation is treated as a thread of execution that can dynamically offload work, such as thinking or retrieving information, by spawning new threads. Thread enables a model to adapt, through recursive spawning, the amount of intermediate work it uses to produce different parts of its token sequence. We apply this framework in the settings of LLM task completion and question answering. We show that Thread, implemented using a few-shot learning approach, achieves state-of-the-art performance on diverse benchmarks in these settings.
Appendix A Illustration of intermediate work used to produce tokens with Thread
@@
Figure 3: Thread allows model to adapt amount of supplemental work used to produce tokens.
Appendix B Implementing Thread for LLM task completion and question answering
We release all code and data at https://github.com/philipmit/thread. As described in section  3, we implement Thread with join synchronization for the problems evaluated in this paper. We implement a spawning mechanism using a special stop token, $\omega_{\_}{listen}$. A thread ends when it generates the end token, $\omega_{\_}{end}$. We use => as $\omega_{\_}{listen}$ and END as $\omega_{\_}{end}$.
@@ Appendix B Implementing Thread for LLM task completion and question answering
As shown in Algorithm 2 below, threads can get feedback from the environment the same way they get feedback from a child thread. They generate tokens that represent the action and then produce the token $\omega_{\_}{listen}$ to listen for feedback from the environment. As shown in Algorithm 2, the action, $a$, is executed in the environment, $E$, and the output, $o=E(a)$, is appended to the thread’s token sequence before it continues generation. The function $\xi$ parses the action from the token sequence. As with prior methods such as ReAct [37], actions are specified following the ‘>’ token.
@@
To ensure consistency across benchmarks and with prior work, we use Meta-Llama-3-8B for Llama-3-8b, Llama-2-7b-hf for Llama-2-7b, and CodeLlama-7b-hf for CodeLlama-7b available on Huggingface [35] and use gpt-4-0613 for GPT-4 and gpt-3.5-turbo-instruct for GPT-3.5 from the OpenAI API, with the temperature set to 0 for all experiments.
@@
Table 8 shows the mean and standard error of the performance of Thread with GPT-3.5 across 5 runs on each benchmark. Overall, the results are consistent across multiple runs.
Table 8: Mean and standard error of Thread performance with GPT-3.5.
Benchmark Mean (Std. Err.)
DataCommons QA 76.7 (.17)
MIMIC-III ICU QA 71.4 (.36)
ALFWorld 95.7 (.28)
TextCraft 93.7 (.37)
WebShop 48.6 (.40)
Appendix D Ablations with parent-child organization and exchange of information
For the implementation of Thread for the settings evaluated in this paper (described in section  3), to improve the ability of the parent and child threads to efficiently organize, update, and exchange shared information, we allow the model to define variables representing pieces of information that are important to the given task. Figure 4 shows the results with and without using these variables for GPT-3.5 (left) and Llama-3-8b (right). Overall, we see that the performance increases slightly with GPT-3.5 and more significantly with Llama-3-8b in the implementation that allows the model to create variables. Nonetheless, the performance of both implementations show a significant improvement over prior methods with the same models.
Figure 4: Performance of Thread for implementations with (dark blue) and without (gray) allowing model to create variables for GPT-3.5 (left) and Llama-3-8b (right).
Appendix E DataCommons QA and MIMIC-III ICU QA Benchmarks

### v2

@@ Results.
Thread outperforms prior methods by at least 20% absolute points with GPT-3.5 and at least 40% points with Llama-3-8b and 30% points CodeLlama-7b (Table 3). Further, Llama-3-8b with Thread outperforms GPT-3.5 with all prior methods and CodeLlama-7b with Thread outperforms GPT-3.5 with all prior methods except TDAG.
4.3 WebShop
WebShop is an online shopping environment with over a million real-world products Yao et al. (2022). The benchmark requires the model to interact with the website to purchase a product based on specifications provided by a user. The evaluation metrics include the success rate, which is the percentage of products that were purchased with full success, and the score, which is the average percentage of desired attributes covered by the purchased items. Following prior work Shinn et al. (2024); Prasad et al. (2023); Zhou et al. (2023), we evaluate Thread on a test set of 100 instructions.
Table 3: TextCraft success rate (%). *As reported in Prasad et al. (2023).
@@ 4.3 WebShop
Reflexion* Shinn et al. (2024) 32.0 - - -
ReAct Yao et al. (2023) 20.5 12.5 8.0 10.5
DecomP Khot et al. (2023) 68.5 47.0 12.0 38.5
ADaPT Prasad et al. (2023) 52.5 23.5 12.0 18.0
TDAG Wang et al. (2024) 73.5 48.5 14.0 31.0
@@
Requires
ext. mem.
Method GPT-3.5 Llama-3-8b Llama-2-7b CodeLlama-7b
@@
Yes Reflexion Shinn et al. (2024) 38 64.4 32 59.8 8 19.7 17 57.3
LATS Zhou et al. (2023) 40 76.0 34 61.5 12 30.1 21 60.7
RAP* Kagaya et al. (2024) 48 76.1 - - - - - -
AutoGuide* Fu et al. (2024) 46 73.4 - - - - - -
No ReAct Yao et al. (2023) 37 59.5 31 54.1 10 18.5 17 49.2
DecomP Khot et al. (2023) 43 58.7 35 56.3 14 29.6 21 57.1
ADaPT Prasad et al. (2023) 44 60.0 35 58.2 13 28.7 21 56.4
TDAG Wang et al. (2024) 45 64.5 37 63.5 14 29.8 23 58.5
@@ Results.
Table 4 shows the success rate and score for each method. We again separate the results based on whether the method requires the model to have access to external memory. With GPT-3.5, Thread outperforms other prompt-only methods by an absolute 4% in success rate and over 10% in score and outperforms RAP Kagaya et al. (2024) by 1% in success rate (with a similar score). Thread achieves an absolute 10%, or greater, improvement in success rate with Llama-3-8b and CodeLlama-7b. Llama-3-8b with Thread achieves a higher success rate than GPT-3.5 with all prior methods, with the exception of RAP.
@@ 4.4 DataCommons QA
DataCommons QA is a benchmark consisting of questions that can be answered using data provided by Google DataCommons Guha (2019) https://datacommons.org. The questions range from comparing statistics in different locations to making predictions regarding future trends. The test set includes a total of 140 questions. We implement Thread with a few-shot prompt and compare its performance to three baselines: Reflexion Shinn et al. (2024), Natural Language Embedded Programs (NLEP) Zhang et al. (2023), and NLEP+ReAct, an extension of NLEP that allows the model to evaluate intermediate outputs of its analysis as it answers the question.
@@ Results.
Table 5 shows the accuracy of each method on DataCommons QA. Thread outperforms prior methods by over 10% absolute points with GPT-3.5, Llama-3-8b, and CodeLlama-7b. Llama-3-8b with Thread outperforms GPT-3.5 with all prior methods.
@@
Reflexion Shinn et al. (2024) 37.9 24.3 10.7 20.7
DecomP Khot et al. (2023) 64.3 57.9 15.7 31.4
NLEP Zhang et al. (2023) 28.6 21.4 11.4 19.3
@@
Reflexion Shinn et al. (2024) 38.1 19.4 8.8 16.9
DecomP Khot et al. (2023) 61.9 51.3 13.1 30.6
NLEP Zhang et al. (2023) 35.6 17.5 8.1 15.6
@@
4.5 MIMIC-III ICU QA
The MIMIC-III ICU QA benchmark consists of patient-focused questions based on clinical time-series data made available by MIMIC-III Johnson et al. (2016). The benchmark reflects a setting in which a healthcare provider can ask natural language questions about patients in the intensive care unit (ICU) and receive an answer from the language model based on the relevant patient data. The test set includes 160 questions. Similar to above, we implement Thread with a few-shot prompt and compare its performance relative to Reflexion Shinn et al. (2024), NLEP, and NLEP+ReAct.
Results.
As shown in Table 6, Thread significantly outperforms prior methods across all models except Llama-2-7b. All methods show the lowest performance with Llama-2-7b, which is consistent across all benchmarks.
@@ 5 Related Work
The Thread framework relates to prior approaches for dynamically adapting work, or computational steps, used during model generation. Previous approaches for dynamically adapting computation based on the given problem require special training Goyal et al. (2023); Nye et al. (2022) or novel model architectures Graves (2016); Banino et al. (2021); Dehghani et al. (2018). Thread provides a more general and flexible framework that, as we show, can be applied without modifications to the underlying model architecture and without further training. In addition, unlike methods that allow the model to adapt the amount of internal computations Goyal et al. (2023); Graves (2016); Banino et al. (2021); Dehghani et al. (2018), the intermediate work used to supplement model generation with Thread can involve work that extends beyond internal thinking, such as retrieving information or interacting with an external environment. In addition, Thread allows for greater interpretability, since all of the intermediate work is performed by generating meaningful tokens.
@@
In this paper, we apply Thread in a setting where it improves the ability of LLMs to decompose a given problem into simpler sub-problems. There has been significant prior work involving the decomposition of problems with neural modular modeling architectures Andreas et al. (2016); Talmor and Berant (2018); Min et al. (2019); Jiang and Bansal (2019); Gupta et al. (2019); Perez et al. (2020); Khot et al. (2021). Later work has used fine-tuning or in-context learning with modern LLMs for decomposition with multi-step tasks Khot et al. (2023); Wang et al. (2023), mathematical reasoning Gao et al. (2023); Lee and Kim (2023), and program synthesis Murali et al. (2018); Nye et al. (2019); Zheng et al. (2023). To handle complex tasks, a more recent approach, called ADaPT, has improved upon these methods by using a planner to further decompose a task when a failure is encountered Prasad et al. (2023), which was further extended to a multi-agent setting with TDAG Wang et al. (2024). Unlike prior methods, such as Decomposed Prompting Khot et al. (2023), ADaPT and TDAG enable the model to adapt a plan upon task failure. In contrast to these approaches, Thread has the advantage of allowing the model to adapt its decision-making in real time as it receives feedback from each step. With Thread, each step of the plan is specified only after receiving feedback from the previous step. In addition, Thread enables the model to flexibly specify and, if needed, re-specify sub-tasks when a sub-task fails or when unexpected feedback is returned. Finally, Thread, which uses the same prompting for every thread at every step of the problem, provides a more unified framework than ADaPT and TDAG, which involve separate prompting and few-shot examples for the “planner” and “executor” modules. With these methods, the planner module outlines a full plan and the executor follows all steps of the plan. If the executor fails, the planner is tasked with further decomposing the failed step into more sub-steps. Instead of outlining a full plan and attempting all steps, Thread involves specifying one step, allocating this to a child thread, and receiving feedback from the child thread before proceeding.
@@ 6 Conclusion
We introduce Thread, a general framework in which model generation is treated as a thread of execution that can dynamically offload work, such as thinking or retrieving information, by spawning new threads. Thread enables a model to adapt, through recursive spawning, the amount of intermediate work it uses to produce different parts of its token sequence. We apply this framework in the settings of agent tasks and question answering. We show that Thread, implemented using a few-shot learning approach, achieves state-of-the-art performance on diverse benchmarks.
Acknowledgements
This work was supported by the MIT-IBM Watson AI Lab and Quanta Computer, Inc. We also thank the Google DataCommons team for their help with developing the DataCommons QA benchmark.
Limitations
The implementation of Thread in this paper does not involve any explicit mechanisms for error handling and instead relies on the inherent reasoning ability of LLMs to respond to unexpected feedback from child threads or from the environment. An advantage of Thread is that individual threads can fail without directly affecting other threads. However, the context necessary to self-correct may be lost if error reporting is not handled appropriately. More work is needed to develop robust error detection and recovery mechanisms to ensure that the information necessary for self-correction is preserved and utilized effectively. In addition, our implementation of Thread involves limited communication between the parent and child threads, where the context for the child thread is based on the last line of the parent’s token sequence. This can result in the child missing information that could help with its work, leading to less efficient or effective generation. Overall, more work is need to improve the propagation of information, as defined by the functions $\phi$ and $\psi$, between parent and child threads.
Ethical Statement
In this work, we show how Thread can be applied to improve LLM agent task completion and question answering. LLMs are being increasingly deployed to autonomously interact with external environments and humans. By improving this capacity, our work has the potential for amplifying risk associated with automated decision-making or facilitate harmful use of LLMs. Addressing these risks requires careful consideration of ethical guidelines, robustness checks, and transparency in algorithm deployment to ensure that advancements in LLMs contribute positively to societal welfare.
Appendix A Illustration of Intermediate Work used to Produce Tokens
@@
Appendix B Implementing Thread for Agent Tasks and Question Answering
We release all code and data at https://github.com/philipmit/thread. As described in section  3, we implement Thread with join synchronization for the problems evaluated in this paper. We implement a spawning mechanism using a special stop token, $\omega_{listen}$. A thread ends when it generates the end token, $\omega_{end}$. We use => as $\omega_{listen}$ and END as $\omega_{end}$.
@@ Appendix B Implementing Thread for Agent Tasks and Question Answering
As shown in Algorithm 2 below, threads can get feedback from the environment the same way they get feedback from a child thread. They generate tokens that represent the action and then produce the token $\omega_{listen}$ to listen for feedback from the environment. As shown in Algorithm 2, the action, $a$, is executed in the environment, $E$, and the output, $o=E(a)$, is appended to the thread’s token sequence before it continues generation. The function $\xi$ parses the action from the token sequence. As with prior methods such as ReAct Yao et al. (2023), actions are specified following the ‘>’ token.
@@
To ensure consistency across benchmarks and with prior work, we use Meta-Llama-3-8B for Llama-3-8b, Llama-2-7b-hf for Llama-2-7b, and CodeLlama-7b-hf for CodeLlama-7b available on Huggingface Wolf et al. (2019) and use gpt-4-0613 for GPT-4 and gpt-3.5-turbo-instruct for GPT-3.5 from the OpenAI API, with the temperature set to 0 for all experiments.
Table 7: Mean and standard error of Thread performance and thread depths with GPT-3.5.
Benchmark Mean (Std. Err.) Avg. thread depth Max thread depth
DataCommons QA 76.7 (.17) 4.1 6
MIMIC-III ICU QA 71.4 (.36) 4.3 6
ALFWorld 95.7 (.28) 3.7 7
TextCraft 93.7 (.37) 5.8 10
WebShop 48.6 (.40) 3.6 7
Model All Pick Clean Heat Cool Look Pick2
GPT-3.5 95.5 100 96.8 82.6 95.2 100 100
Llama-3-8b 49.3 58.3 38.7 56.5 57.1 72.2 11.8
CodeLlama-7b 61.9 41.7 87.1 65.2 61.9 38.9 64.7
@@
Table 7 shows the mean and standard error of the performance of Thread with GPT-3.5 across 5 runs on each benchmark. Overall, the results are consistent across multiple runs. Table 7 also shows the average and max thread depths for each task.
Appendix D Task-general Prompting for ALFWorld
Table 8 shows the results when testing Thread with task-general prompting in ALFWorld. To implement the task-general prompt, we split the prompt into one set of examples for the main thread and a second set of examples for all other threads. The same sets of examples used for all tasks. GPT-3.5 achieves the same combined success rate with the task-general prompting as it does with the task-specific prompting. Further, while the performance of Llama-3-8b and CodeLlama-7b degrades with task-general prompting, it remains a significant improvement over task-specific prompting with prior methods (Table 2).
Appendix E Error and Ablation Analysis
E.1 Error Types
To identify what types of errors Thread reduces, we classify the errors that are responsible for each method’s failures. Descriptions and examples of each error type are provided in Tables 9 and 10 for the agent tasks and Tables 11 and 12 for the QA tasks. We classified these errors through hand review of each failure case, where the first error type to occur within the failure case was identified as the error type responsible for the failure. Figures 4a, 4b, 5, and 6a show the error counts across different methods and tasks. We focus this analysis on the prompt-only methods to provide a more clear evaluation of how the novel aspects of Thread change performance relative to other methods that leverage few-shot learning.
E.2 Ablations
Figures 4c, 4d, and 6b show how the error counts change when applying the following three modifications to Thread:
  1. 1.
    Removing the variables used by parent and child threads to manipulate and organize shared information.
  2. 2.
    Replacing $\phi$ with a function that returns the parent thread’s full token sequence.
  3. 3.
    Using a separate prompt for Thread 0 (which performs the initial task decomposition) instead of using the same few-shot prompt for all threads.
E.3 Agent Tasks
Task decomposition and flexible sub-task specification reduce failures caused by inadequate plans. In Figures 4a and 4b, we see that the methods that involve task decomposition (i.e., Decomposed prompting, ADaPT, and Thread) show fewer failures that are caused by inadequate planning. This is likely because there is a dedicated line of reasoning for generating the plan, instead of it being part of the single line of model generation that is responsible for both defining the plan and carrying out the steps of the plan, like in ReAct.
We see that Thread, which allows for flexible sub-task specification, further reduces failures caused by inadequate planning. This is due to the fact that, when a child thread fails, the parent thread can accommodate the child thread’s feedback, adapt the sub-task specification, and spawn a new thread. Prior methods do not allow for this flexible sub-task re-specification, as they either require the model to select from a fixed set of pre-defined sub-task handlers (e.g., Decomposed Prompting) or do not accommodate sub-task feedback when handling failures (e.g, ADaPT and TDAG). Figure 7 shows an example where a child thread, which is tasked with checking a cabinet for a plate, returns feedback saying the cabinet is closed. This is a common problem, where a child thread (or a sub-task handler/executor in Decomposed Prompting or ADaPT) return some unpredictable intermediate result instead of fully completing its task (which, in this case, involves opening the cabinet and checking for a plate). To address this with Thread, the parent thread spawns a new child thread with a modified sub-task specification (which includes opening the cabinet and checking if there is a plate) in order to prevent the new child thread from repeating the mistake made by the previous child. Prior methods do not allow for these dynamic, nuanced adjustments to the plan to accommodate the full range of possible intermediate feedback involved in complex tasks.
Shortening action sequences reduces failures caused by misinterpreting environment feedback. We see that errors in interpreting environment feedback are lowest with Decomposed Prompting and Thread. One potential explanation for this is that, by decomposing the task into more sub-components than other methods, Decomposed Prompting and Thread require each line of model generation to execute shorter sequences of consecutive actions. As a result, each line of model generation needs to manage shorter action-observation sequences and, therefore, is able to more accurately interpret and keep track of the information returned from the environment.
Figure 5 shows the number of times GPT-3.5 misinterprets environment feedback when executing action sequences of different lengths when using the different methods. By significantly reducing the number of cases where the model has to perform long action sequences, Decomposed Prompting and Thread reduce the number of interpretation errors.
E.4 QA Tasks
Task decomposition reduces failures that are caused by performing the wrong analysis. In Figure 6a, we see that Decomposed Prompting and Thread show fewer failures caused by the model performing the wrong analysis for the given question. This is similar to what we see with the agent tasks where problem decomposition improves the planning. By separating the line of model generation that defines the plan (defining the analysis type) from the line of model generation that is responsible for executing the plan (completing the analysis), the model can more effectively accomplish both tasks.
Flexible sub-task specification reduces failures that are caused by runtime errors. We see that Thread shows the fewest failures that are caused by runtime errors. With Thread, when an error occurs, the error is returned to the parent thread, which can adapt its plan by modifying the sub-task specification and spawning a new thread (as discussed above). In addition, based on the ablation analysis (Figure 6b), we see that the variables used in Thread may also reduce failures due to runtime errors, especially with smaller models such as Llama-3-8b.
Figure 4: Failure counts in ALFWorld and TextCraft for different methods (a and b) and for modified versions of Thread (c and d).
[…]
Figure 6: Failure counts, combined for DataCommons QA and MIMIC-III ICU QA, for different methods (a) and modified versions of Thread (b).
Table 9: Description of error types for agent tasks.
Error type Description
Inadequate plan The plan does not include steps sufficient to complete the task (including steps to recover from unexpected intermediate results).
Invalid action The model attempts to perform an invalid action in the environment.
Misinterprets environment feedback The model incorrectly interprets information returned from the environment.
Table 10: Examples of error types for agent tasks.
Error type Examples
Inadequate plan ALFWorld: The plan does not include cleaning the plate when the model is tasked with putting a clean plate on the table. The plan does not involve opening the cabinet to check for a plate when an unexpected sub-task result is returned (e.g., “The cabinet is closed”). TextCraft: The plan involves crafting an object before acquiring one of the precursors. The plan does not finish checking if a material (e.g., bamboo) can be retrieved from the environment when an unexpected result is returned (e.g., “Could not find valid recipe for bamboo.”).
Invalid action ALFWorld: The model incorrectly specifies the action for examining an object with the desklamp. TextCraft: The model incorrectly states a crafting recipe.
Misinterprets environment feedback ALFWorld: The model continues looking for a plate after the environment indicates “You see a plate 1”. TextCraft: The model proceeds as though it has successfully fetched bamboo despite the environment returning “Could not find bamboo”.
Other ALFWorld: The model reaches the maximum number of actions allowed in the ALFWorld environment. TextCraft: The model runs out of context.
Table 11: Description of error types for QA tasks.
Error type Description
Retrieves wrong information The model does not retrieve the information needed to answer the question.
Performs wrong analysis The model retrieves the correct information, but does not perform the analysis needed to answer the question.
Misinterprets results of analysis The model performs the correct analysis, but misinterprets the results.
Runtime error Error occurs during the execution of the model’s code.
Table 12: Examples of error types for QA tasks.
Error type Examples
Retrieves wrong information The model retrieves data on smoking prevalence of all people when asked specifically about the female population.
Performs wrong analysis The model compares values from 2014 to 2021 when asked to compare values from 2018 to 2021.
Misinterprets results of analysis The model indicates a variable is increasing despite its analysis showing that it is decreasing.
Runtime error The model tries to concatenate an integer to a string.
Other The model runs out of context.
Figure 7: Example of task in ALFWorld where ADaPT fails to productively respond to the feedback from a sub-task, while Thread successfully accommodates the feedback to re-specify a new sub-task.
Appendix F DataCommons QA and MIMIC-III ICU QA Benchmarks